OpenStudy (anonymous):

Simplify the given expression. square root of negative 10 end root times square root of negative 8 negative 4 square root 5 4 square root of 5 4 I square root of 5 negative 4 I square root of 5

OpenStudy (anonymous):

\[\sqrt{-10} * \sqrt{-8}\]

OpenStudy (anonymous):

start with\[\sqrt{10}i\times \sqrt{8}i\]

OpenStudy (anonymous):

then since \(i^2=-1\) you have \[-\sqrt{80}\] but you still have to write \(\sqrt{80}\) in simplest radical form sort of like the other ones

OpenStudy (anonymous):

you got that part?

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

i dont know where to start to write square root of 80 in rad form

OpenStudy (anonymous):

you need to find a perfect square inside 80

OpenStudy (anonymous):

16 x 5 = 80

OpenStudy (anonymous):

just like you wrote \(27=9\times 3\) where \(9\) was the perfect square

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

\[4\sqrt{5}\]

OpenStudy (anonymous):

yay! don't forget the minus sign out front

OpenStudy (anonymous):

\[-4\sqrt5\]

OpenStudy (anonymous):

\[-4\sqrt{5}\] is the answer?

OpenStudy (anonymous):

oui